Year 2014 2015 2016 2017 2018

Dividend $0.8 $0.9 $1.00 $1.3 $1.4

$ Change ? ?   ? ?

% Change     ? ?   ? ?

Given the dividend information above, if the dividend just paid is $1.4 in 2018 and the price per share of this company’s stock is $64, what is the cost of equity using the dividend growth model?
Answer #1

Change % = (1.4/0.8)1/4 -1 = 15.02%

Calculating Cost of Equity,

Cost of Equity = 1.4(1.1502)/(64) + 0.1502

Cost of Equity = 17.54%